OpenStudy (anonymous):

i give medal, and become fan What is the approximate radius of a sphere whose volume is 1,349 cm³? (1 point) 5.7 cm 6.9 cm 11 cm 14.7 cm

OpenStudy (anonymous):

4/3 * pi* r^3=1349 r^3= 1349*3/4* 3.14------------r= 6.857=6.9 cm

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

merry christmas